How do you solve the number of gifts given in the 12 Days of Christmas algebraically using a progression formula? The answer is not 78.
Thanks

Answer by fractalier(2101)

The sum of n consecutive integers, starting with one, is
S = [n(n+1) / 2] so that for n = 12, we have
S = (12)(13) / 2 = 78
That IS the answer if you give one on the 1st day, two on the 2nd, etc...